OpenStudy (anonymous):

2. Using the description below, illustrate and label the process that is being completed with all the appropriate lab equipment. a) A student diluted a copper (II) sulphate pentahydrate stock solution by placing a 1 mL of stock in 3 mL of dH2O. b) The sample prepared in step a) was further diluted by placing 0.50 mL of the copper (II) sulphate pentahydrate into 4.5 mL of dH2O. c) Lastly, 3 mL of the sample prepared in step b) was added to 12 mL dH2O. What is the total dilution factor? If one was told that the stock solution

OpenStudy (aaronq):

a) 1:4 or \(\dfrac{1}{4}\) b) 1:10 or \(\dfrac{1}{10}\) c) 1:4 or \(\dfrac{1}{4}\) multiply them.

OpenStudy (anonymous):

How did you get those?

OpenStudy (aaronq):

for example: a) a 1 mL of stock in 3 mL of dH2O. there are 4 mL in total, and your initial was 1 mL so the ratio is \(\dfrac{1}{4}\) if you had 3 mL of stock and added 7 mL of water then you'd have \(\dfrac{3}{10}\), makes sense?